# About testing Bell locality at colliders

### v2

@@ 5. Quantum tomography: a new take on loopholes
Low-energy experiments with photons (or solid-state devices) can be considered more akin to proper Bell tests than those conducted at colliders because they can be formulated as black boxes. [^4] The extent to which these experiments are device-independent must be evaluated on a case-by-case basis, see for example  [51]. In this black-box approach, the settings of the polarimeters serve as inputs, while the measured frequencies act as outputs (for a captivating illustration of this concept, refer to [66]). However, local hidden variable models pose a challenge to these tests by exploiting potential loopholes through which local variables could mimic Bell nonlocality.
@@
The applicability of these loopholes has been extended to encompass Bell nonlocality measurements conducted at colliders. In these discussions, the construction of suitable local hidden variable models is frequently done only in abstract terms. The reason for this is straightforward: such models must simultaneously evade the stringent limits imposed by low-energy tests and bypass the high-energy tests. These requirements make the very definition of local hidden variable models suitable for collider experiments a daunting task.
@@
Regardless of whether all loopholes are closed or not, establishing the violation of Bell inequalities in new settings, such as those provided by collider experiments, is of interest because, at the very least, it makes the construction of local hidden-variable models alternative to quantum mechanics more challenging. We believe there is more to it, though, due to the new features introduced by quantum tomography.
@@
All the loopholes so far discussed in the literature exploit the incomplete experimental determination of the bipartite state that is provided in the traditional setup of a Bell test, notably in the two-photon system. Two directions, arbitrary but fixed, are defined for each of the two polarization measurements performed on each subsystem; correlations among these polarization measurements are then compared and the Bell inequality evaluated. These four directions are the context of the measurement; correlations are computed on some statistical distributions of the variables.
@@
Loopholes bypass nonlocality by encoding information on the measurement directions in additional variables. This way, the correlations entering the Bell inequality are altered once this information is transmitted. This is achieved either through an exchange of information that makes the direction $\vec{n}_{i}$, along which the first particle’s spin is measured, known to the second particle whose spin is measured along the direction $\vec{n}_{j}$—as in the case of the space-time locality loophole—or by using this information to alter the events to be measured—as in the detection loophole. This key observation provides a new perspective on all loopholes affecting Bell tests. It is crucial for the existence of loopholes that the measurement directions are set at some point, otherwise, no information can be used to manipulate correlations and consequently satisfy the inequality. This point is often overlooked in discussions of loopholes in Bell tests, especially when their effects are discussed in abstract terms.
@@
Quantum tomography provides us with the state of the bipartite system under exam, a system that we could then probe to assert the potential Bell inequality violation. However, if the state—which is nonlocal— is given with sufficient precision, there is no way to bypass the mathematics that connects operators, states, and expectation values. The presence of entanglement and the potential violation of Bell inequalities can then be straightforwardly established through the mentioned observables, within the uncertainties proper of the tomographic measurements. There is no room for loopholes to act: loopholes do not affect states, quantum tomography nor quantum mechanics. Loopholes exploit the fact that only partial information is made available in a Bell test where the freedom of manipulating the setting necessarily leaves room for the loopholes to slip in.
@@
In the following we briefly review the different loopholes affecting Bell tests one at a time, showing how the tomographic procedure made available by collider experiments allows to disregard them.
5.1. Space-time locality
@@
At collider experiments, the condition of space-like separation between the polarization measurements requires, at first blush, particles with the same lifetime and mass. Yet this restriction is actually not necessary because the time-like separation between production and measurement points is only an issue for experiments with the traditional settings. At colliders, in fact, no specific setting is decided beforehand as the whole density matrix is fully reconstructed. The directions involved in a Bell test are decided only afterwards, when the reconstructed state is pulled through the machinery of quantum mechanics to compute the required correlation values. There is no way for the state at the production point, nor for the two subsystem composing the final state, to know this choice in advance. For this reason, even for two decays taking place one inside the future cone of the other—as in the case of two particles with very different masses and decay lengths—the loophole cannot be invoked.
5.2. Detection
@@ 5.3. Freedom of choice
${\mathcal{P}}(\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})=\int\mbox{d}\lambda\,p_{\lambda}(\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})\eta(\lambda|\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})\,$, (5.1)
one should take into account possible correlations between the hidden variables and the detector settings [73], which would modify the above relation thereby allowing a hidden variable theory to pass the Bell test. This loopohole essentially proposes a limited form of a super-deterministic model in which, however, only a tiny amount of information needs to be exchanged [74].
This loophole is considered impossible to close through scientific methods in the traditional setup (although see the objection presented in the next Section). In low-energy tests it has been made more and more unlikely by making the choice of detector settings be determined by a large ensemble of pseudo-random events (canvassed from a large sampling of human choices) [75] or events that are in the very (cosmological) far past [76].
At collider experiments, no detector settings are chosen. There is therefore no information that the hidden variables can carry to alter the reconstruction of the state which proceeds always in the same manner. Likewise, the loophole cannot alter the subsequent probing of correlations in the Bell inequality, which is solely regulated by quantum mechanics.
5.4. Free will
There exists a stronger form of the previous loophole that consists in supposing that the hidden variable distribution contains the outcome of the experiment about to be performed. For particle accelerators, this would mean that the hidden variables regulate the angular distribution of the momenta involved in the tomographic procedure. The same variables must also include information about the detector [^6] For instance, about the direction of the magnetic field used in the identification of charged particles. because the contextuality requirement [77].
This possibility appears incompatible with physics because it necessitates an absolute determination of the initial conditions [78], which in turn requires information on all the infinitely many digits comprising the real numbers used to specify these conditions [79]. Furthermore, this vast amount of information would have to be confined within a finite volume of spacetime, thereby violating the Bekenstein bound [80]. It seems to us that this fundamental principle, which states that initial conditions cannot be specified with arbitrary precision, is often overlooked in discussions on determinism. [^7] The burden for those who want to revive super-determinism would be to demonstrate that the number of digits required to fix the initial conditions and enable predictability for a sufficiently long period (e.g., the age of the Universe) is sufficiently small to fall below the Bekenstein bound. Remember though that even if one disregards the above argument and insists on entertaining such a philosophical stance, the contextuality requirement renders full-blown determinist models rather unappealing.
5.5. Memory
@@ 5.6. Coincidence
6. If they exist, hidden variables are new physics
As we contemplate the future of quantum tomography at colliders, two unexpected outcomes can emerge. If the experimental data yield correlations surpassing the Tsirelson bound [62], this result would imply that nonlocality surpasses what is predicted by quantum mechanics. All the tests conducted at colliders have so far confirmed quantum mechanics. Alternatively, the reconstructed state might not be an acceptable state within the Hilbert space of quantum mechanics. This could occur if the experimental values of certain cross sections violate unitarity. In such cases, the problem would be addressed by seeking new physics rather than being interpreted as a signal of the failure of quantum mechanics.
Coming to hidden variables, they were initially proposed for theories that were not fundamental, such as atomic physics. On the other hand, particle physics delves into the most fundamental level of nature, where there is no room for hiding. This leads to speculations about whether the incompleteness of quantum mechanics can be rephrased as the incompleteness of the Standard Model interactions and particle content. While it is possible that a clever hidden variable model could reproduce the states identified by the tomographic procedure, regardless of the model, it will be contextual and nonlocal, unless we consider the super-deterministic choice, which we already ruled out due to the required infinite amount of information. If we attempt to construct such a model (an example of which, though non-relativistic, is Bohm’s model [83]), we will face the challenge of doing so at the microscopic level and therefore by means of new particles or interactions that carry the information encoded in the hidden variables. These features cannot remain hidden in a fundamental theory and must manifest themselves as these new (currently elusive) particles or interactions with special properties. From the perspective of particle physics, the search for hidden variable models can be framed as the search for physics beyond the Standard Model.
7. Outlook
In this paper, we have introduced a Bell inequality tailored to spin correlation matrices constructed at colliders by means of quantum state tomography. It provides a quantum witness to nonlocality. We have demonstrated that realistic particle decays can violate this inequality. These collider tests of Bell inequalities are device-dependent, but we have also argued that they appear to circumvent the typical loopholes encountered in low-energy Bell tests. Furthermore, we have stressed that super-determinism, the most extreme of all loopholes, lacks physical basis and must be rejected.
Quantum tomography has been successfully performed using experimental data from colliders. Specifically, analyses of $B$-meson decays at Belle II and LHCb [84], Charmonium decays at BESIII [85] and pions from kaon decays [86] have been reinterpreted to achieve this. These processes have demonstrated Bell nonlocality with a significance exceeding $5\sigma$ in a manner, as we argued here, that avoids any loopholes. Consequently, these results extend the low-energy findings to encompass weak and strong interactions, as well as qutrits systems. Furthermore, the detection of entanglement in top-quark pair production data at the LHC has been achieved with a significance of $5\sigma$ [87, 88, 89].
Analyses using $\tau$ lepton pairs [90, 91], top-quark pairs [92, 93, 94, 95, 96, 97, 98] and vector gauge bosons [99, 100, 101, 102, 103] show that similar results can be expected in the upcoming years for many other processes. The prospects of testing Bell inequality at future colliders has also been discussed [104, 102, 105, 103, 106, 107].

### v3

@@ Appendix A Quantum tomography: a new take on loopholes
Low-energy experiments with photons (or solid-state devices) can be considered more akin to proper Bell tests than those conducted at colliders because they can be formulated as black boxes. [^7] The extent to which these experiments are device-independent must be evaluated on a case-by-case basis, see for example  [51]. In this black-box approach, the settings of the polarimeters serve as inputs, while the measured frequencies act as outputs (for a captivating illustration of this concept, refer to [103]). However, local hidden variable models pose a challenge to these tests by exploiting potential loopholes through which local variables could mimic Bell inequality violations.
@@
The applicability of these loopholes has been extended to encompass Bell non-locality measurements conducted at colliders. In these discussions, the construction of suitable local hidden variable models is frequently done only in abstract terms. The reason for this is straightforward: such models must simultaneously evade the stringent limits imposed by low-energy tests and bypass the high-energy tests. These requirements make the very definition of local hidden variable models suitable for collider experiments a daunting task.
@@
Regardless of whether all loopholes are closed or not, establishing some form of Bell non-locality in new settings, such as those provided by collider experiments, is of interest because, at the very least, it makes the construction of local hidden-variable models alternative to quantum mechanics more challenging. We believe there is more to it, though, due to the new features introduced by quantum tomography.
@@
Most of the loopholes so far discussed in the literature exploit the incomplete experimental determination of the bipartite state that is provided in the traditional setup of a Bell test, notably in the two-photon system. Two directions, arbitrary but fixed, are defined for each of the two polarization measurements performed on each subsystem; correlations among these polarization measurements are then compared and the Bell inequality evaluated. These four directions are the context of the measurement; correlations are computed on some statistical distributions of the variables.
@@
Loopholes bypass non-locality by encoding information on the measurement directions in additional variables. This way, the correlations entering the Bell inequality are altered once this information is transmitted. This is achieved either through an exchange of information that makes the direction $\vec{n}_{i}$, along which the first particle’s spin is measured, known to the second particle whose spin is measured along the direction $\vec{n}_{j}$—as in the case of the space-time locality loophole—or by using this information to alter the events to be measured—as in the detection loophole. This key observation provides a new perspective on the applicability of loopholes affecting Bell tests. It is crucial for the existence of loopholes that the measurement directions are set at some point, otherwise, no information can be used to manipulate correlations and consequently satisfy the inequality. This point is often overlooked in discussions of loopholes in Bell tests, especially when their effects are discussed in abstract terms.
@@
Quantum tomography provides us with the state of the bipartite system under exam, a system that we could then probe to witness the potential for Bell non-locality. If the state—which is non-local— is given with sufficient precision, there is no way to bypass the mathematics that connects operators, states, and expectation values. The presence of entanglement and Bell non-locality can then be straightforwardly established through the mentioned observables, within the uncertainties proper of the tomographic measurements. There is no room for loopholes to act: loopholes do not affect states, quantum tomography nor quantum mechanics. Most loopholes exploit the fact that only partial information is made available in a Bell test where the freedom of manipulating the setting necessarily leaves room for the loopholes to slip in.
@@
Since in the literature, there exist various attempts to extend the discussion of these loophole in the framework of collider physics, in the following we briefly review the different loopholes affecting Bell tests one at a time, showing how the tomographic procedure made available by collider experiments allows to disregard them.
A.1. Space-time locality
@@
At collider experiments, the condition of space-like separation between the polarization measurements requires, at first blush, particles with the same lifetime and mass. Yet this restriction is actually not necessary because the time-like separation between production and measurement points is only an issue for experiments with the traditional settings. At colliders, in fact, no specific setting is decided beforehand as the whole density matrix is fully reconstructed. The directions involved in witnessing Bell non-locality are decided only afterwards, when the reconstructed state is pulled through the machinery of quantum mechanics to compute the required correlation values. There is no way for the state at the production point, nor for the two subsystem composing the final state, to know this choice in advance. For this reason, even for two decays taking place one inside the future cone of the other—as in the case of two particles with very different masses and decay lengths—the loophole cannot be invoked.
A.2. Detection
@@ A.3. Freedom of choice
${\mathcal{P}}(\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})=\int\mbox{d}\lambda\,p_{\lambda}(\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})\eta(\lambda|\uparrow_{\hat{n}_{i}};\uparrow_{\hat{n}_{j}})\,$, (A.1)
one should take into account possible correlations between the hidden variables and the detector settings [73], which would modify the above relation thereby allowing a hidden variable theory to pass the Bell test. This loophole essentially proposes a limited form of a super-deterministic model in which, however, only a tiny amount of information needs to be exchanged [110].
This loophole is considered impossible to close through scientific methods in the traditional setup (although see the objection presented in the main text). In low-energy tests it has been made more and more unlikely by making the choice of detector settings be determined by a large ensemble of pseudo-random events (canvassed from a large sampling of human choices) [111] or events that are in the very (cosmological) far past [112].
At collider experiments, no detector settings are chosen. There is therefore no information that the hidden variables can carry to alter the reconstruction of the state which proceeds always in the same manner. Likewise, the loophole cannot alter the subsequent probing of correlations and the witnessing of Bell non-locality, which are solely regulated by quantum mechanics.
A.4. Memory